OpenStudy (cucacula):

Determine the nature of the roots of the following quadratic equation. x² + 4x +4 = 0 anyone know how to answer this?

OpenStudy (anonymous):

Ok, you know the quadratic formula?

OpenStudy (cucacula):

what is the Quadratic formula? @freewilly922

OpenStudy (anonymous):

there are two basic natures of roots they are probably looking for. Real and imaginary roots. Imaginary roots result from a square root of a negative number somewhere. \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

a good word to google is "discriminant"

OpenStudy (anonymous):

So the quadratic formula will give you the roots for any quadratic equation... the only way to get a complex root is for the square root part to be taking the square root of a negative number.

OpenStudy (cucacula):

The answer is 2 ?? @freewilly922

OpenStudy (anonymous):

That would be the number of roots, but that's not what was asked. The question asked about the nature of the roots. Whether they are REAL or COMPLEX I assume.

OpenStudy (anonymous):

So Real roots would have \[\sqrt{\text{positive#}}\] and Complex Roots would have \[\sqrt{\text{negative#}}\]

OpenStudy (anonymous):

For example if the quadratic formula gave you \[\frac{-(9)\pm\sqrt{(9)^2-4(10)(10)}}{(2)(10)}\] You would end up with the square root of -319 in the fraction and that would give you complex roots. \[\frac{-(9)\pm\sqrt{(9)^2-4(10)(1)}}{(2)(10)}\] on the other hand would give you the square root of 41 and that would leave you with real roots. As was mentioned earlier, $\[b^2-4ac\] is generally called the disciminant and when it is negative, you have complex roots and when positive you have Real roots.

OpenStudy (cucacula):

so (9)² - 4(10)(1) = 41 which b² - 4ac > 0 right? @freewilly922

OpenStudy (anonymous):

Yes for my example. That is correct:) so that one has two -real- roots.

OpenStudy (anonymous):

Can you now do this with the problem they gave you?

OpenStudy (anonymous):

\[b^2 -4ac > 0: \text{2 real roots}\] \[b^2 -4ac < 0: \text{2 complex roots}\] \[b^2 -4ac = 0: \text{1 real root*}\] * I think there is a special name for this root. Like it is degree 2 or something.

OpenStudy (cucacula):

so the asnwer is real and equal :) Thanks @freewilly922